# On De Giorgi’s lemma for variational interpolants

### v2

@@ 4 The Banach case: results
In the upcoming Section 4.1 we will show that the De Giorgi’s estimate previously proved in [MRS13a, Lemma 6.1] in the radially differentiable case, in fact improves to an equality. Then, in Sec. 4.3 we will drop the radial differentiabiilty condition and extend the validity of De Giorgi’s estimate to general dissipation potentials.
4.1 Equality in De Giorgi’s estimate for radially differentiable potentials
@@
$\Phi_{\sigma}(u^{o};u)=\sigma\,{\mathcal{R}}\big{(}\frac{1}{\sigma}(u{-}u^{o})\big{)}+{\mathcal{E}}(u).$
@@
In particular, the left and right derivatives of $\sigma\mapsto\Phi_{\sigma}(u^{o};u)$ exist, see Lemma 3.12, and are independent of the energy ${\mathcal{E}}$. Moreover, the one-sided derivatives are ordered such that ${\mathrm{D}}^{-}_{\sigma}\Phi(u^{o};\sigma,w)\leq{\mathrm{D}}^{+}_{\sigma}\Phi(u^{o};\sigma,w)<0$ for every $w\in J_{\sigma}(u^{o})$.
@@
However, because of the supremum over ${\mathrm{D}}^{-}_{\sigma}\Phi$ and the infimum over ${\mathrm{D}}^{+}_{\sigma}\Phi$, in general one cannot expect to generate one chain of inequalities from the two estimates in (3.25). Even if $J_{\sigma}(u^{o})$ is single-valued, one still has the wrong estimate.
@@
The only case where the marginal estimates (3.25) are useful, is exactly when ${\mathcal{R}}$ is radially differentiable, cf. (3.16). Then ${\mathrm{D}}^{-}_{\sigma}\Phi={\mathrm{D}}^{+}_{\sigma}\Phi$ implies $\delta^{-}_{\sigma}(u^{o})\geq\delta^{+}_{\sigma}(u^{o})$ and (3.25) leads to
$\liminf_{h\to 0^{+}}\frac{1}{h}\big{(}\phi(u^{o};\sigma)-\phi(u^{o};\sigma{-}h)\big{)}\geq\delta^{-}_{\sigma}(u^{o})\geq\delta^{+}_{\sigma}(u^{o})\geq\limsup_{h\to 0^{+}}\frac{1}{h}\big{(}\phi(u^{o};\sigma{+}h)-\phi(u^{o};\sigma)\big{)}.$
From this, De Giorgi’s identity follows easily, and a fortiori (4.1b) below for any measurable selection $\widetilde{\xi}_{\sigma}\in\partial_{\mathcal{R}}{\mathcal{E}}(u^{o},\sigma;\widetilde{u}_{\sigma})$. The existence of such selections will be addressed in the upcoming Lemma 4.11.
Theorem 4.1 (De Giorgi’s identity with radial differentiability)
Consider the
$\mathrm{gBGS}$ $(X,{\mathcal{E}},{\mathcal{R}})$ safisfying Hypothesis 3.1. Moreover, assume that ${\mathcal{R}}$ is radially differentiable, i.e. (3.17) or equivalently (3.16) holds. Fix $u^{o}\in\mathrm{dom}({\mathcal{E}})$.
@@
Then, every measurable variational interpolant $(0,\infty)\ni\sigma\mapsto\widetilde{u}_{\sigma}\in X$ fulfills De Giorgi’s identity
${\mathcal{E}}(\widetilde{u}_{\sigma})+\sigma{\mathcal{R}}\big{(}\frac{1}{\sigma}(\widetilde{u}_{\sigma}{-}u^{o})\big{)}+\int_{0}^{\sigma}\mathcal{C}_{\mathcal{R}}(u^{o},\rho;\widetilde{u}_{\rho})\;\!\mathrm{d}\rho={\mathcal{E}}(u^{o})\,,$ (4.1a)
and in particular there holds
@@
$\forall\,\sigma>0:\quad{\mathcal{E}}(\widetilde{u}_{\sigma})+\sigma{\mathcal{R}}\big{(}\frac{1}{\sigma}(\widetilde{u}_{\sigma}{-}u^{o})\big{)}+\int_{0}^{\sigma}{\mathcal{R}}^{*}(-\widetilde{\xi}_{\rho})\;\!\mathrm{d}\rho={\mathcal{E}}(u^{o})$ (4.1b)
for every measurable selection $\sigma\mapsto\widetilde{\xi}_{\sigma}\in\partial_{\mathcal{R}}{\mathcal{E}}(u^{o},\sigma;\widetilde{u}_{\sigma})$.
@@
Proof. We fix a small $r>0$ and consider the marginal function $[r,\sigma]\ni\rho\mapsto\phi(u^{o};\rho)$. According to Proposition 3.13 $\phi$ is Lipschitz and hence differentiable almost everywhere in $[r,\sigma]$. This implies that $\phi^{\prime}_{-}$ and $\phi^{\prime}_{+}$ exist and coincide almost everywhere.
@@
$\phi^{\prime}(\sigma)=-{\mathcal{R}}^{*}(\eta)\qquad\text{for all }\eta\in\partial{\mathcal{R}}\big{(}\frac{1}{\sigma}(\widetilde{u}_{\sigma}{-}u^{o})\big{)}\,.$
@@
$\phi^{\prime}(\sigma)=-\mathcal{C}_{\mathcal{R}}(u^{o},\sigma;\widetilde{u}_{\sigma})\quad\text{for a.a.}\,\sigma\in(0,\infty)\,.$
@@
$\phi(u^{o};\sigma)={\mathcal{E}}(\widetilde{u}_{\sigma})+\sigma{\mathcal{R}}\big{(}\frac{1}{\sigma}(\widetilde{u}_{\sigma}{-}u^{o})\big{)}=\phi(u^{o};r)-\int_{r}^{\sigma}\mathcal{C}_{\mathcal{R}}(u^{o},\rho;\widetilde{u}_{\rho})\;\!\mathrm{d}\rho\,.$
The superlinearity of ${\mathcal{R}}$ implies $\phi(u^{o};r)\to{\mathcal{E}}(u^{o})$ as $r\to 0^{+}$, see (3.5). Therefore, taking the limit $r\to 0^{+}$ gives the desired De Giorgi’s identity (4.1).
We conclude this section with an example in which ${\mathcal{R}}$ is not radially differentiable but De Giorgi’s identity still holds.
@@
For $u^{o}=6$ we obtain the unique variational interpolant
@@
It can easily be checked that identity (3.13) holds for all $\sigma>0$, where the choice $\widetilde{\xi}_{\sigma}={\mathrm{D}}{\mathcal{E}}(\widetilde{u}_{\sigma})=\widetilde{u}_{\sigma}$ is mandatory because the Fréchet subdifferential of ${\mathcal{E}}$ is single-valued. For $\sigma\in[2,5]$ we have $-\widetilde{\xi}(\sigma)=-6{+}\sigma\in\partial{\mathcal{R}}(-1)=[-4,-1]$, where we used $\frac{1}{\sigma}(\widetilde{u}_{\sigma}{-}6)=-1$.
@@ 4.2 Regularity of the variational interpolant
Indeed, on the one hand, we will require uniform convexity of the mapping $u\mapsto\Phi_{\sigma}(u^{o};\cdot)$, a sufficient condition for which is, of course, uniform convexity of the energy functional ${\mathcal{E}}$. On the other hand, we will need to reinforce radial differentiability. Indeed, while the latter property is equivalent to the fact that the composed, a priori multi-valued, mapping ${\mathcal{R}}^{*}{\circ}\partial{\mathcal{R}}:X\rightrightarrows X^{*}$ is single-valued (cf. Proposition 3.8), we will now further require that ${\mathcal{R}}^{*}{\circ}\partial{\mathcal{R}}$ is Lipschitz continuous (on bounded subsets of $X$).
@@
    There exist $\overline{\lambda}>0$ and $\sigma_{*}>0$ such that for all $\sigma>0$ the mapping $u\mapsto\Phi_{\sigma}(u^{o};u)$ is $\overline{\lambda}$-convex, namely
    $\exists\overline{\lambda}>0\ \ \forall\,\sigma>0\,,\ \forall\,u_{0},u_{1}\in X\,,\ \forall\,\theta\in[0,1]\,:$ (4.2)
    $\Phi_{\sigma}(u^{o};(1{-}\theta)u_{0}{+}\theta u_{1})\leq(1{-}\theta)\Phi_{\sigma}(u^{o};u_{0})+\theta\Phi_{\sigma}(u^{o};u_{1})-\frac{\overline{\lambda}}{2}\theta(1{-}\theta)\|u_{0}{-}u_{1}\|^{2}\,.$
@@
    The mapping ${\mathcal{R}}^{*}{\circ}\partial{\mathcal{R}}:X\rightrightarrows X^{*}$ (locally) Lipschitz continuous:
    $\forall\,M>0\ \exists\,C_{M}>0\ \forall\,(u_{1},\xi_{1}),\,(u_{2},\xi_{2})\in X{\times}X^{*},\ \max_{i=1,2}\|u_{i}\|\leq M,\ \xi_{i}\in\partial{\mathcal{R}}(u_{i})\,:$ (4.3)
@@
Before stating our result, let us pin down two key consequences of Hypothesis 4.3:
  • It follows from (4.2) (cf., e.g., [MRS13b, Prop. 2.4]) that the Fréchet subdifferential $\partial\Phi_{\sigma}(u^{o};\cdot):X\rightrightarrows X^{*}$ can be characterized in terms of the following global estimate: for all $\sigma>0$ and for every $u\in\mathrm{dom}(\partial\Phi_{\sigma}(u^{o};\cdot))$ we have that
    $\begin{gathered}\omega\in\partial\Phi_{\sigma}(u^{o};u)\text{ if and only if }\\ \Phi_{\sigma}(u^{o};v)-\Phi_{\sigma}(u^{o};u)\geq\langle\omega,v{-}u\rangle+\frac{\overline{\lambda}}{2}\|v{-}u\|^{2}\quad\text{for all }v\in X\,.\end{gathered}$ (4.4)
  • By Prop. 3.8, (4.3) in particular implies that for every $M>0$ the restriction of ${\mathcal{R}}$ to the ball $\overline{B}_{M}(0)$ is radially differentiable. Therefore, recalling Lemma 3.9 we have that for every $v\in\overline{B}_{M}(0)$ the function
    $\begin{gathered}g(v;\rho)=\rho{\mathcal{R}}\left(\frac{1}{\rho}v\right)\text{ is differentiable at every }\rho>0,\text{ with }\\ \frac{\;\!\mathrm{d}}{\;\!\mathrm{d}\rho}g(v;\rho)={\mathcal{R}}\left(\frac{1}{\rho}v\right)-\langle\eta,\frac{1}{\rho}v\rangle={\mathcal{R}}^{*}(\eta)\qquad\text{for all }\eta\in\partial{\mathcal{R}}\left(\frac{1}{\rho}v\right)\,.\end{gathered}$ (4.5)
@@
Let the $\mathrm{gBGS}$ $(X,{\mathcal{E}},{\mathcal{R}})$ satisfy Hypotheses 3.1 and 4.3; and let $u^{o}\in\mathrm{dom}({\mathcal{E}})$ be fixed. Then, for all measurable selection $(0,\infty)\ni\sigma\mapsto\widetilde{u}_{\sigma}\in J_{\sigma}(u^{o})$ we have $\widetilde{u}\in\mathrm{Lip}_{\mathrm{loc}}(]0,\infty[;X)$.
Proof. Preliminarily, we observe that, because of $\sup_{\sigma>0}\sigma{\mathcal{R}}((\widetilde{u}_{\sigma}{-}u^{o})/\sigma)\leq C$, we have
@@
Let us fix $\sigma_{\#}>0$ and let $[\sigma_{1},\sigma_{2}]\subset[\sigma_{\#},\infty)$ be arbitrary. We apply (4.4) with $\sigma=\sigma_{1}$, $u=\widetilde{u}_{\sigma_{1}}$, $v=\widetilde{u}_{\sigma_{2}}$ and $\omega=0$ as, indeed, $0\in\partial\Phi_{\sigma_{1}}(u^{o};\widetilde{u}_{\sigma_{1}})$ is the Euler-Lagrange equation for $\widetilde{u}_{\sigma_{1}}\in J_{\sigma_{1}}(u^{o})$. Thus, we obtain
$\frac{\overline{\lambda}}{2}\|\widetilde{u}_{\sigma_{2}}{-}\widetilde{u}_{\sigma_{1}}\|^{2}$ $\leq\Phi_{\sigma_{1}}(u^{o};\widetilde{u}_{\sigma_{2}})-\Phi_{\sigma_{1}}(u^{o};\widetilde{u}_{\sigma_{1}})$
$={\mathcal{E}}(\widetilde{u}_{\sigma_{2}})-{\mathcal{E}}(\widetilde{u}_{\sigma_{1}})+\sigma_{1}{\mathcal{R}}\left(\frac{1}{\sigma_{1}}(\widetilde{u}_{\sigma_{2}}{-}u^{o})\right)-\sigma_{1}{\mathcal{R}}\left(\frac{1}{\sigma_{1}}(\widetilde{u}_{\sigma_{1}}{-}u^{o})\right)\,.$
@@
$\leq\sigma_{2}{\mathcal{R}}\big{(}\frac{1}{\sigma_{2}}(\widetilde{u}_{\sigma_{1}}{-}u^{o})\big{)}-\sigma_{1}{\mathcal{R}}\big{(}\frac{1}{\sigma_{1}}(\widetilde{u}_{\sigma_{1}}{-}u^{o})\big{)}+\sigma_{1}{\mathcal{R}}\big{(}\frac{1}{\sigma_{1}}(\widetilde{u}_{\sigma_{2}}{-}u^{o})\big{)}-\sigma_{2}{\mathcal{R}}\big{(}\frac{1}{\sigma_{2}}(\widetilde{u}_{\sigma_{2}}{-}u^{o})\big{)}$
@@
hence $\sigma\mapsto\widetilde{u}_{\sigma}$ is Lipschitz continuous on $[\sigma_{\#},\infty)$. By the arbitrariness of $\sigma_{\#}$, we conclude $\widetilde{u}\in\mathrm{Lip}_{\mathrm{loc}}(]0,\infty[;X)$.
@@
We conclude this section by gaining further insight into property (4.3). The key observation is that it can be reformulated in terms of the mapping $\mathfrak{P}$ from (3.18).

### v3

@@ 4 The Banach case: results
In the upcoming Section 4.1 we will show that the De Giorgi estimate previously proved in [MRS13, Lemma 6.1] in the radially differentiable case, in fact improves to an equality. Section 4.2 will address the regularity of the variational interpolant. Then, in Sec. 4.3 we will drop the radial differentiabiilty condition and extend the validity of the De Giorgi estimate to general dissipation potentials.
4.1 Equality in the De Giorgi estimate for radially differentiable potentials
@@
$\Phi_{\sigma}(u)=\sigma\,{\mathcal{R}}\big(\frac{1}{\sigma}(u{-}u^{\circ})\big)+{\mathcal{E}}(u).$
@@
In particular, the left and right derivatives of $\sigma\mapsto\Phi_{\sigma}(u)$ exist, see Lemma 3.16, and are independent of the energy $\mathcal{E}$}. Moreover, the one-sided derivatives are ordered such that ${\mathrm{D}}^{-}_{\sigma}\Phi(\sigma,w)\leq{\mathrm{D}}^{+}_{\sigma}\Phi(\sigma,w)<0$ for every $w\in J_{\sigma}$.
@@
However, because of the supremum over ${\mathrm{D}}^{-}_{\sigma}\Phi$ and the infimum over ${\mathrm{D}}^{+}_{\sigma}\Phi$, in general one cannot expect to generate one chain of inequalities from the two estimates in (3.28). Even if $J_{\sigma}$ is single-valued, one still arrived at the wrong estimate.
@@
The only case where the marginal estimates (3.28) are useful to prove the De Giorgi identity, is exactly when $\mathcal{R}$} is radially differentiable, cf. (3.19). Then ${\mathrm{D}}^{-}_{\sigma}\Phi={\mathrm{D}}^{+}_{\sigma}\Phi$ implies $\delta^{-}_{\sigma}\geq\delta^{+}_{\sigma}$ and (3.28) leads to
$\liminf_{h\to 0^{+}}\frac{1}{h}\big(\phi(\sigma)-\phi(\sigma{-}h)\big)\geq\delta^{-}_{\sigma}\geq\delta^{+}_{\sigma}\geq\limsup_{h\to 0^{+}}\frac{1}{h}\big(\phi(\sigma{+}h)-\phi(\sigma)\big).$
From this, the De Giorgi identity follows easily.
We will also address the validity of the identity involving a measurable selection $\widetilde{\xi}_{\sigma}\in\partial_{{\mathcal{R}}}{\mathcal{E}}(\sigma;\widetilde{u}_{\sigma})$ (we will often refer to $\widetilde{\xi}$ as a force selection). For the existence of such a selection (see Lemma 4.3), it is useful to impose the following closedness condition for the conditioned subdifferential $\partial_{{\mathcal{R}}}{\mathcal{E}}:(0,\infty){\times}X\rightrightarrows X^{*}$.
Hypothesis 4.1 (Closedness of $\partial_{{\mathcal{R}}}{\mathcal{E}}$)
The conditioned subdifferential $\partial_{{\mathcal{R}}}{\mathcal{E}}:(0,\infty){\times}X$ $\rightrightarrows X^{*}$ is closed on energy sublevels, i.e.
$\hskip-14.22636pt\forall\,E>0:\ \left\{\begin{array}[]{c}(\sigma_{n},u_{n},\xi_{n})\rightharpoonup(\sigma,u,\xi)\ \text{ in }(0,\infty){\times}X{\times}X^{*},\\ u_{n}\in S_{E}\text{ and }\xi_{n}\in\partial_{{\mathcal{R}}}{\mathcal{E}}(\sigma_{n};u_{n})\text{ for all }n\in{\mathbb{N}}\end{array}\right\}\ \Longrightarrow\ \xi\in\partial_{{\mathcal{R}}}{\mathcal{E}}(\sigma;u)\,.$ (4.3)
Theorem 4.2 (The De Giorgi identity with radial differentiability)
@@
Consider the $\mathrm{gBGS}$ $(X,{\mathcal{E}},{\mathcal{R}})$ satisfying Hypothesis 3.1. Fix $u^{\circ}\in\mathrm{dom}({\mathcal{E}})$ and assume that $\mathcal{R}$} is radially differentiable, i.e. (3.20) or equivalently (3.19) holds. Then, every measurable variational interpolant $(0,\infty)\ni\sigma\mapsto\widetilde{u}_{\sigma}\in J_{\sigma}$ fulfills the De Giorgi identity
${\mathcal{E}}(\widetilde{u}_{\sigma})+\sigma{\mathcal{R}}\big(\frac{1}{\sigma}(\widetilde{u}_{\sigma}{-}u^{\circ})\big)+\int_{0}^{\sigma}\mathcal{C}_{{\mathcal{R}}}(\rho;\widetilde{u}_{\rho})\;\!\mathrm{d}\rho={\mathcal{E}}(u^{\circ})\,.$ (4.4a)
If additionally Hypothesis 4.1 holds, then there exists a measurable force selection $\sigma\mapsto\widetilde{\xi}_{\sigma}\in\partial_{{\mathcal{R}}}{\mathcal{E}}(\sigma;\widetilde{u}_{\sigma})$ such that the improved De Giorgi identity with force selection holds:
@@
$\forall\,\sigma>0:\quad{\mathcal{E}}(\widetilde{u}_{\sigma})+\sigma{\mathcal{R}}\big(\frac{1}{\sigma}(\widetilde{u}_{\sigma}{-}u^{\circ})\big)+\int_{0}^{\sigma}{\mathcal{R}}^{*}(-\widetilde{\xi}_{\rho})\;\!\mathrm{d}\rho={\mathcal{E}}(u^{\circ}).$ (4.4b)
@@
Proof. We fix a small $r>0$ and consider the marginal function $r,\sigma]\ni\rho\mapsto\phi(\rho)$. According to Proposition 3.17 $\phi$ is Lipschitz and hence differentiable almost everywhere in $r,\sigma]$. This implies that $\phi^{\prime}_{-}$ and $\phi^{\prime}_{+}$ exist and coincide almost everywhere.
@@
$\phi^{\prime}(\sigma)=-{\mathcal{R}}^{*}(\eta)\qquad\text{for all }\eta\in\partial{\mathcal{R}}\big(\frac{1}{\sigma}(\widetilde{u}_{\sigma}{-}u^{\circ})\big)\,.$
@@
$\phi^{\prime}(\sigma)=-\mathcal{C}_{{\mathcal{R}}}(\sigma;\widetilde{u}_{\sigma})\quad\text{for a.a.}\,\sigma\in(0,\infty)\,.$
@@
$\phi(\sigma)={\mathcal{E}}(\widetilde{u}_{\sigma})+\sigma{\mathcal{R}}\big(\frac{1}{\sigma}(\widetilde{u}_{\sigma}{-}u^{\circ})\big)=\phi(u^{\circ};r)-\int_{r}^{\sigma}\mathcal{C}_{{\mathcal{R}}}(\rho;\widetilde{u}_{\rho})\;\!\mathrm{d}\rho\,.$
The superlinearity of $\mathcal{R}$} implies $\phi(r)\to{\mathcal{E}}(u^{\circ})$ as $r\to 0^{+}$, see (3.6). Therefore, taking the limit $r\to 0^{+}$ gives the desired De Giorgi identity (4.4a).
The last part of the assertion immediately follows from the upcoming Lemma 4.3.
The following result collects two straightforward consequences of Hypothesis 4.1.
Lemma 4.3
Let the $\mathrm{gBGS}$ $(X,{\mathcal{E}},{\mathcal{R}})$ satisfy Hypotheses 3.1 and 4.1. Then,
  1. 1.
    $\mathcal{C}_{{\mathcal{R}}}(\sigma;u):=\min\big\{\,{\mathcal{R}}^{*}({-}\xi)\>\big|\>\xi\in\partial_{{\mathcal{R}}}{\mathcal{E}}(\sigma;u)\,\big\}\quad\text{for all }u\in J_{\sigma};$ (4.5)
  2. 2.
    For every measurable selection $(0,\infty)\ni\sigma\mapsto\widetilde{u}_{\sigma}\in J_{\sigma}$ lying in some energy sublevel $S_{E}$, there exists a measurable selection
    $\sigma\mapsto\widetilde{\xi}_{\sigma}\in\mathfrak{A}_{{\mathcal{R}}}(\sigma,\widetilde{u}_{\sigma}):=\mathrm{argmin}\big\{\,{\mathcal{R}}^{*}({-}\xi)\>\big|\>\xi\in\partial_{{\mathcal{R}}}{\mathcal{E}}(\sigma;\widetilde{u}_{\sigma})\,\big\}$ (4.6a)
    such that
    $\mathcal{C}_{{\mathcal{R}}}(\sigma;{\widetilde{u}_{\sigma}})={\mathcal{R}}^{*}({-}\widetilde{\xi}_{\sigma})\quad\text{for all }\sigma>0\,.$ (4.6b)
Proof. Property (4.5) is easily checked via the direct method by relying on (4.3), with a similar argument as for (3.11a). For (4.6b), observe that, as a consequence of Hypothesis 4.1, the multivalued mapping
$\mathfrak{A}_{{\mathcal{R}}}:(0,\infty){\times}X\rightrightarrows X^{*};\quad\mathfrak{A}_{{\mathcal{R}}}(\sigma,u):=\mathrm{argmin}\big\{\,{\mathcal{R}}^{*}({-}\xi)\>\big|\>\xi\in\partial_{{\mathcal{R}}}{\mathcal{E}}(\sigma;u)\,\big\}$ (4.7)
is upper semicontinuous with respect to convergence in $\mathbb{R}$} for $\sigma$ and weak convergence for $u$. Therefore, for every measurable selection $(0,\infty)\ni\sigma\mapsto\widetilde{u}_{\sigma}\in J_{\sigma}$ the multivalued mapping $\sigma\mapsto\mathfrak{A}_{{\mathcal{R}}}(\sigma,\widetilde{u}_{\sigma})$ is measurable, as it is given by the composition of measurable mappings. Then, [CaV77, Thm. 3.22] grants the existence of a measurable selection $\sigma\mapsto\widetilde{\xi}_{\sigma}\in\mathfrak{A}_{{\mathcal{R}}}(\sigma,\widetilde{u}_{\sigma})$.
We conclude this section with an example in which $\mathcal{R}$} is not radially differentiable but the De Giorgi identity still holds.
@@
For $u^{\circ}=6$ we obtain the unique variational interpolant
@@
It can easily be checked that identity (3.16) holds for all $\sigma>0$, where the choice $\widetilde{\xi}_{\sigma}={\mathrm{D}}{\mathcal{E}}(\widetilde{u}_{\sigma})=\widetilde{u}_{\sigma}$ is mandatory because the Fréchet subdifferential of $\mathcal{E}$} is single-valued. For $\sigma\in[2,5]$ we have $\widetilde{\xi}(\sigma)=-6{+}\sigma\in\partial{\mathcal{R}}(-1)=[-4,-1]$, where we used $\frac{1}{\sigma}(\widetilde{u}_{\sigma}{-}6)=-1$.
Recalling the definitions $\delta^{\pm}_{\sigma}$ from (3.28) we obtain the strict inequalities $\delta^{-}_{\sigma}=-{\mathcal{R}}^{*}(-4)=-\frac{7}{2}\lneqq\partial_{\sigma}\phi(\sigma)=\sigma{-}\frac{11}{2}\lneqq-\frac{1}{2}=-{\mathcal{R}}^{*}(-1)=\delta^{+}_{\sigma}$ for $\sigma\in(2,5)$.
@@ 4.2 Regularity of the variational interpolant
Furthermore, on the one hand, we will require uniform convexity of the mapping $u\mapsto\Phi_{\sigma}(u)$, a sufficient condition for which is, of course, uniform convexity of the energy functional $\mathcal{E}$}. On the other hand, we will need to reinforce radial differentiability. Indeed, while the latter property is equivalent to the fact that the composed, a priori multivalued, mapping ${\mathcal{R}}^{*}{\circ}\partial{\mathcal{R}}:X\rightrightarrows X^{*}$ is single-valued (cf. Proposition 3.11), we will now further require that $\mathcal{R}$$^{*}{\circ}\partial{\mathcal{R}}$ is Lipschitz continuous (on bounded subsets of $X$).
@@
    There exist $\overline{\lambda}>0$ and $\sigma_{*}>0$ such that for all $\sigma>0$ the mapping $u\mapsto\Phi_{\sigma}(u)$ is $\overline{\lambda}$-convex, namely
    $\hskip-14.22636pt\exists\,\overline{\lambda}>0\ \ \forall\,\sigma>0\,,\ \forall\,u_{0},u_{1}\in X\,,\ \forall\,\theta\in[0,1]\,:$ (4.8)
    $\hskip-14.22636pt\Phi_{\sigma}((1{-}\theta)u_{0}{+}\theta u_{1})\leq(1{-}\theta)\Phi_{\sigma}(u_{0})+\theta\Phi_{\sigma}(u_{1})-\frac{\overline{\lambda}}{2}\theta(1{-}\theta)\|u_{0}{-}u_{1}\|^{2}\,.$
@@
    The mapping ${\mathcal{R}}^{*}{\circ}\partial{\mathcal{R}}:X\rightrightarrows X^{*}$ is Lipschitz continuous on bounded sets, i.e.,
    $\hskip-14.22636pt\forall\,M>0\ \exists\,C_{M}>0\ \forall\,(v_{1},\xi_{1}),\,(v_{2},\xi_{2})\in X{\times}X^{*},\ \max_{i=1,2}\|v_{i}\|\leq M,\ \xi_{i}\in\partial{\mathcal{R}}(v_{i})\,:$ (4.9)
@@
Before stating our result, let us pin down two key consequences of Hypothesis 4.5:
  • It follows from (4.8) (cf., e.g., [MRS13, Prop. 2.4]) that the Fréchet subdifferential $\partial\Phi_{\sigma}(\cdot):X\rightrightarrows X^{*}$ can be characterized in terms of the following global estimate: for all $\sigma>0$ and for every $u\in\mathop{\mathrm{dom}}(\partial\Phi_{\sigma}(\cdot))$ we have that
    $\begin{gathered}\omega\in\partial\Phi_{\sigma}(u)\text{ if and only if }\\ \Phi_{\sigma}(v)-\Phi_{\sigma}(u)\geq\langle\omega,v{-}u\rangle+\frac{\overline{\lambda}}{2}\|v{-}u\|^{2}\quad\text{for all }v\in X\,.\end{gathered}$ (4.10)
  • By Prop. 3.11, (4.9) in particular implies that for every $M>0$ the restriction of $\mathcal{R}$} to the ball $\overline{B}_{M}(0)$ is radially differentiable. Therefore, recalling Lemma 3.12 we have that for every $v\in\overline{B}_{M}(0)$ the function
[…]
@@
Let the $\mathrm{gBGS}$ $(X,{\mathcal{E}},{\mathcal{R}})$ satisfy Hypotheses 3.1 and 4.5; and let $u^{\circ}\in\mathrm{dom}({\mathcal{E}})$ be fixed. Then, for all measurable selection $(0,\infty)\ni\sigma\mapsto\widetilde{u}_{\sigma}\in J_{\sigma}$ we have $\widetilde{u}\in\mathrm{Lip}_{\mathrm{loc}}(]0,\infty[;X)$.
Proof. Preliminarily, we observe that, because of $\sup_{\sigma>0}\sigma{\mathcal{R}}((\widetilde{u}_{\sigma}{-}u^{\circ})/\sigma)\leq C$, we have
@@
Let us fix $\sigma_{\#}>0$ and let $\sigma_{1},\sigma_{2}]\subset[\sigma_{\#},\infty)$ be arbitrary. We apply (4.10) with $\sigma=\sigma_{1}$, $u=\widetilde{u}_{\sigma_{1}}$, $v=\widetilde{u}_{\sigma_{2}}$ and $\omega=0$ as, indeed, $0\in\partial\Phi_{\sigma_{1}}(\widetilde{u}_{\sigma_{1}})$ is the Euler-Lagrange equation for $\widetilde{u}_{\sigma_{1}}\in J_{\sigma_{1}}(u^{\circ})$. Thus, we obtain
$\frac{\overline{\lambda}}{2}\|\widetilde{u}_{\sigma_{2}}{-}\widetilde{u}_{\sigma_{1}}\|^{2}$ $\leq\Phi_{\sigma_{1}}(\widetilde{u}_{\sigma_{2}})-\Phi_{\sigma_{1}}(\widetilde{u}_{\sigma_{1}})$
$={\mathcal{E}}(\widetilde{u}_{\sigma_{2}})-{\mathcal{E}}(\widetilde{u}_{\sigma_{1}})+\sigma_{1}{\mathcal{R}}\left(\frac{1}{\sigma_{1}}(\widetilde{u}_{\sigma_{2}}{-}u^{\circ})\right)-\sigma_{1}{\mathcal{R}}\left(\frac{1}{\sigma_{1}}(\widetilde{u}_{\sigma_{1}}{-}u^{\circ})\right)\,.$
@@
$\leq\sigma_{2}{\mathcal{R}}\big(\frac{1}{\sigma_{2}}(\widetilde{u}_{\sigma_{1}}{-}u^{\circ})\big)-\sigma_{1}{\mathcal{R}}\big(\frac{1}{\sigma_{1}}(\widetilde{u}_{\sigma_{1}}{-}u^{\circ})\big)+\sigma_{1}{\mathcal{R}}\big(\frac{1}{\sigma_{1}}(\widetilde{u}_{\sigma_{2}}{-}u^{\circ})\big)-\sigma_{2}{\mathcal{R}}\big(\frac{1}{\sigma_{2}}(\widetilde{u}_{\sigma_{2}}{-}u^{\circ})\big)$
@@
hence $\sigma\mapsto\widetilde{u}_{\sigma}$ is Lipschitz continuous on $\sigma_{\#},\infty)$. By the arbitrariness of $\sigma_{\#}$, we conclude $\widetilde{u}\in\mathrm{Lip}_{\mathrm{loc}}(]0,\infty[;X)$.
@@
We conclude this section by gaining further insight into property (4.9). The key observation is that it is implied by radial differentiability and superlinearity of $\mathcal{R}$$^{*}$.